OpenStudy (anonymous):

please help me five years ago the ratio of alice's to bernards age was 1:4. 10 years from now, the ratio of alices to bernards age will be 2:3. how old is bernard? a)15 b)17 c)25 d)27 e)30

OpenStudy (anonymous):

@aum can you please help me with this one too

OpenStudy (anonymous):

This is a system of two linear equations. Our first equation is \[\frac {Alice - 5}{Bernard - 5} = \frac {1}{4}\] and our second equation is \[\frac {Alice + 10}{Bernard + 10} = \frac {2}{3}\] Solving these equations will give us the ages of both Alice and Bernard.